1. Compute the following limits. 9x4 - 4y4 (a) lim lim (x,y)–(0,0) 3x2 + 2y2
Answer #1

9x4 – 4y4 lim (x,y)=(0,0) 3x2 + 2y2 We can rewrite it as 9x4 – 4y4 lim (x,y)=(0,0) 3x2 + 2y2 = (3x2 + 2y2)(3x2 – 2y2) lim (x,